# AD-LLM: Benchmarking Large Language Models for Anomaly Detection

### v1

@@ 5.2 Problem Statement and Designs
Given a dataset $\mathcal{D}=\{x_{1},x_{2},\dots,x_{n}\}$ and a set of AD models =$\mathcal{M}=\{M_{1},M_{2},\dots,M_{m}\}$, the task is to identify a suitable model $M^{*}\in\mathcal{M}$ using a pre-trained LLM $f_{\text{LLM}}$, based solely on provided information about the dataset and the candidate models.
@@
Figure 4: Comparison of AD performance across five datasets. “LLM Recommendation” denotes the average of models selected by querying the LLM five times (duplicates allowed). “Best Performance” indicates the highest performance among all AD models for each dataset, while “Baseline Average” shows the mean performance of all baseline AD models (as if choosing it randomly). Results show that models recommended by the LLM generally outperform the baseline average and, in some cases, approach the best-performing model, showing the potential of LLM-based UMS.
@@ 5.3 Results, Insights, and Future Directions
For the UMS scenario, which requires sophisticated reasoning, we use GPT-o1-preview OpenAI (2024c) with enhanced reasoning abilities.
LLM recommendations demonstrate strong potential. Figure 4 compares the detection performance across five datasets using three references: (i) the best achievable result from any model, representing the performance upper bound; (ii) the average performance of all baseline models, simulating a scenario where a user selects models at random; and (iii) the results from the LLM-recommended models (ours). The LLM-recommended models surpass the baseline average in most cases, and sometimes approach the best-performing model (e.g., on “IMDB Reviews” and “SMS Spam”). This indicates that LLMs can identify promising AD models using only public information, w/o reliance on historical performance or domain specialists.
LLM justifications lack dataset-specific insights. Although the LLM selects models that perform relatively well, its explanations often remain generic and do not clearly link model selection to specific dataset characteristics. For example, in the “AG News” dataset, the LLM alternated between recommending “OpenAI + LUNAR” and “OpenAI + ECOD,” justifying choices with broad statements like “effective for high-dimensional data” or “parameter-free scalability.” Such non-specific rationales diminish interpretability and user trust, especially when understanding the rationale behind model choice is important.
Future Direction 5: Refine Input Specificity and Addressing Biases. Future work should explore how to provide more dataset-specific details and mitigate potential LLM biases. Ambiguous or incomplete input information may cause the LLM to favor well-known models or those frequently encountered during training. Ensuring detailed and balanced inputs, and exploring how inherent biases in LLMs affect recommendations, will be important steps to improve the fairness and reliability of LLM-based UMS Dai et al. (2024).
@@
6 Conclusion and Future Directions
In this work, we presented AD-LLM, the first comprehensive benchmark that integrates LLMs into three core aspects of anomaly detection in NLP: detection, data augmentation, and model selection. Our results show that LLMs hold strong potential across these tasks. For detection, LLMs demonstrate effective zero-shot and few-shot capabilities, often performing competitively against traditional methods without task-specific training. For data augmentation, both synthetic data generation and category description augmentation lead to improved performance for unsupervised and LLM-based AD methods. For model selection, LLMs can recommend suitable AD models using only general dataset and model descriptions. These findings highlight the emerging role of LLMs in making AD more flexible, data-efficient, and adaptable.
Future Directions. Investigating methods to systematically integrate external context into prompts can reduce variability and enhance precision. Techniques like quantization or hardware optimizations can lower the computational burden for real-world applications. Specialized augmentation approaches, designed to optimize user-prioritized metrics, may further improve AD performance. Finally, improving the specificity and clarity of LLM justifications can enhance interpretability and user trust, and expanding the AD-LLM benchmark to include additional tasks and applications in different fields Huang et al. (2024b); Li et al. (2024b).
@@ 6 Conclusion and Future Directions
AD-LLM explores the application of LLMs in enhancing AD through zero-shot detection, data augmentation, and model selection. These contributions have the potential to significantly improve real-world AD systems in critical areas such as healthcare, finance, and cybersecurity. By enabling robust, adaptable, and efficient solutions for AD tasks, this research empowers practitioners to deploy systems that are responsive to novel challenges while reducing reliance on labeled data and extensive domain expertise.
@@
This study adheres to ethical guidelines, emphasizing considerations around fairness, transparency, and privacy in developing and applying LLM-based AD systems. We emphasize the importance of evaluating and mitigating biases in LLM recommendations, ensuring that outputs are equitable and unbiased. Moreover, privacy is preserved by relying on public data and avoiding the collection of sensitive information.
@@
Despite promising results, several limitations remain. First, our evaluation is constrained to a narrow set of datasets with clear normal-anomaly distinctions, and our settings in AD and category descriptions in DA follow the structure of these datasets, limiting applicability to various domains with ambiguous anomaly definitions. Second, the synthetic data generation in DA is limited in scale, producing few samples per category, and challenges in scaling up remain unresolved. Third, UMS depends on simplistic input data and matching mechanisms. Furthermore, biases in LLM recommendations, such as favoring well-documented or familiar models, need further investigation. Finally, our study evaluates only a subset of popular LLMs, lacking a comprehensive assessment. Additionally, we do not explore few-shot learning or fine-tuning, which are widely adopted techniques for enhancing LLM performance and could offer valuable complementary insights for AD tasks.
@@ B.1 Prompt Details
  • Explanation and Implicit CoT. We require an explanation $r$ generated before the anomaly score $s$ for each inquiry as shown in Eq. (1). When generating the explanation, LLMs implicitly create the CoT in the background Liu et al. (2024c). This approach aligns with the auto-regressive nature of decoder-only LLMs, encouraging them to think carefully and logically before determining the anomaly score, thereby enhancing reliability.
In our experiments, we discovered that Llama 3.1 requires implicit CoT. Presenting the anomaly score $s$ before the explanation $r$ causes the Llama 3.1-based detector to crash and consistently outputs $s=0$. This issue does not impact GPT-4o-based detectors. We attribute this to GPT-4o’s significantly larger parameter count (8B), which grants it a stronger resilience to prompt changes.
@@ B.2 Complete Baseline Results
We present a bar chart in Figure 2 comparing the precision and recall LLM-based AD between two settings across five datasets. The detailed results, including precise numerical values, are provided in Table A3 for reference.
Table A3: LLM-based AD precision-recall comparison between two settings across five datasets. The anomaly score threshold $=0.5$ is used to decide the anomaly. The better results for each detector are highlighted in bold.
Settings AG News BBC News IMDB Reviews N24News SMS Spam
Precision $\uparrow$ Recall $\uparrow$ Precision $\uparrow$ Recall $\uparrow$ Precision $\uparrow$ Recall $\uparrow$ Precision $\uparrow$ Recall $\uparrow$ Precision $\uparrow$ Recall $\uparrow$
Llama 3.1 8B Instruct
(1) with $\mathcal{C}_{\text{normal}}$ 0.5364 0.5645 0.5319 0.4098 0.2946 0.7328 0.1130 0.6053 0.1352 0.8961
(2) with $\mathcal{C}_{\text{normal}}$, $\mathcal{C}_{\text{anomaly}}$ 0.2719 0.9518 0.3355 0.8226 0.3379 0.9622 0.2930 0.9067 0.1816 1.000
GPT-4o
(1) with $\mathcal{C}_{\text{normal}}$ 0.5673 0.9058 0.8966 0.8387 0.6739 0.8130 0.7077 0.3206 0.4495 0.6364
(2) with $\mathcal{C}_{\text{normal}}$, $\mathcal{C}_{\text{anomaly}}$ 0.5527 0.9259 0.9048 0.9194 0.6803 0.9428 0.8828 0.9234 0.5891 0.9870
@@ B.3 Error Analysis
In our experiments, LLMs occasionally encountered errors. Llama 3.1 sometimes entered infinite loops, repeatedly generating the same sentence until reaching the token limit, as illustrated in the AG News example shown in Text Box B.3, where the repetitive nonsensical response is highlighted in italics. However, the underlying causes of such behavior remain underexplored. For GPT-4o, outputs were sometimes restricted by OpenAI’s safety filters when addressing potentially harmful content, such as politically sensitive topics.
We exclude these errors from our results but provide the error counts in both “Normal Only” and “Normal + Anomaly” settings for reference in Table A4 for Llama 3.1 and Table A5 for GPT-4o. Notably, the error counts vary between the two settings, suggesting that the triggers for errors, such as Llama’s infinite loop or GPT-4o’s safety filters, are sensitive to prompt variations. This occurs even though the prompts in both settings have similar semantic meanings.
@@
Table A6: Comparison of AD performance across five datasets. “LLM Recommendation” denotes the average of models selected by querying the LLM five times (duplicates allowed). “Best Performance” indicates the highest performance among all AD models for each dataset, while “Baseline Average” shows the mean performance of all baseline AD models (as if choosing it randomly).
Training Set AG News BBC News IMDB Reviews N24 News SMS Spam
LLM Recommendation 0.8908 0.6193 0.6992 0.1573 0.6652 0.2608 0.7706 0.4047 0.5774 0.1548
@@ C.1.1 Evaluation Protocol and Prompt Details
As discussed in §4.2, we set a scenario with limited training data $\mathcal{D}_{\text{small\_train}}=\{x_{1},x_{2},\dots,x_{m}\}$. Specifically, $\mathcal{D}_{\text{small\_train}}$ contains $v$ samples for each normal category $\mathcal{C}_{\text{normal}}^{j}\in\mathcal{C}_{\text{normal}}=\{\mathcal{C}_{\text{normal}}^{1},\dots,\mathcal{C}_{\text{normal}}^{k}\}$, where $k$ is the number of normal categories.
We employ a multi-step strategy in this task to mitigate repetitive outputs, token limit constraints, and difficulties in handling long contexts. The detailed pipeline is outlined below:
@@
    Keywords Generation. To ensure a consistent synthetic data distribution compared with the original training data, $t$ groups of keywords are generated for each normal category $\mathcal{C}_{\text{normal}}^{j}$. We construct the prompt $\mathcal{P}_{\text{keywords}}$ using a template $T_{\text{keywords}}(\cdot)$ as shown in Table  A11. This template utilizes {name} and {original_task} information from Table  A.1. The prompt $\mathcal{P}_{\text{keywords}}$ is processed by the LLM $f_{\text{LLM}}(\cdot)$ to produce $t\times k$ groups of keywords $\mathcal{K}=\{\mathcal{K}_{1},\mathcal{K}_{2},\dots,\mathcal{K}_{t\times k}\}$. Each keyword group $\mathcal{K}_{i}$ contains three keywords with increasing levels of granularity from coarse to fine.
@@
    Synthetic Sample Generation. We iterate the groups of keywords, constructing $\mathcal{P}_{\text{synth}}=\{\mathcal{P}_{\text{synth}}^{1},\mathcal{P}_{\text{synth}}^{2},\dots,\mathcal{P}_{\text{synth}}^{t\times k}\}$ using a template $T_{\text{synth}}(\cdot)$ as displayed in Table  A12. Each prompt $\mathcal{P}_{\text{synth}}^{j}$ is fed into the LLM $f_{\text{LLM}}(\cdot)$ to generate a corresponding synthetic sample $\hat{x}_{j}$. Finally, we obtain a synthetic dataset $\mathcal{D}_{\text{synth}}=\{\tilde{x}_{1},\tilde{x}_{2},\dots,\tilde{x}_{t\times k}\}$.
@@
[…]
$\mathcal{K}=f_{\text{LLM}}(\mathcal{P}_{\text{keywords}})=\{\mathcal{K}_{1},\dots,\mathcal{K}_{t\times k}\}$
$\mathcal{P}_{\text{synth}}=\{T_{\text{synth}}(\mathcal{K}_{1}),\dots,T_{\text{synth}}(\mathcal{K}_{t\times k})\}$
$\mathcal{D}_{\text{synth}}=\{f_{\text{LLM}}(\mathcal{P}_{\text{synth}}^{1}),\dots,f_{\text{LLM}}(\mathcal{P}_{\text{synth}}^{t\times k})\}$
@@ C.1.2 Experiments Details and Challenges
We set the number of samples from each normal category $\mathcal{C}_{\text{normal}}^{j}$ in the limited training set $\mathcal{D}_{\text{small\_train}}$ to $v=10$. Similarly, the number of synthetic samples generated for each normal category $\mathcal{C}_{\text{normal}}^{j}$ in the synthetic set $\mathcal{D}_{\text{synth}}$ is set to $t=50$ for the “AG New”, “BBC News”, “IMDB Reviews”, and “SMS Spam” datasets. For the “N24 News” dataset, we set $v=3$ and $t=12$ to account for its numerous normal categories.
@@
In our experiments, GPT-4o was used for synthetic data generation. We observed that increasing $t$ occasionally caused GPT-4o to terminate the keyword generation process before reaching the token limit. A similar issue occurred with Llama 3.1, even for smaller values of $t$. As a result, Llama 3.1 was excluded from this task. We presume these issues stem from the inherent challenges LLMs face in processing long contexts.
A straightforward solution is generating keywords in multiple rounds with different temperatures and prompt templates. Further techniques, such as sample-wise decomposition and dataset-wise decomposition Long et al. (2024), can be explored to optimize the synthetic data generation.

### v3

@@ 5.2 Problem Statement and Designs
Given a dataset $\mathcal{D}=\left\{x_{1},x_{2},\dots,x_{n}\right\}$ and a set of AD models =$\mathcal{M}=\left\{M_{1},M_{2},\dots,M_{m}\right\}$, the task is to identify a suitable model $M^{*}\in\mathcal{M}$ using a pre-trained LLM $f_{\text{LLM}}$, based solely on provided information about the dataset and the candidate models.
@@
Table 3: The top 2 frequent picks made by each LLM. Counts are aggregated over 25 queries (5 per dataset).
LLMs Top-2 Picks (counts)
OpenAI-o1-preview OpenAI+LUNAR (13), OpenAI+ECOD (8)
DeepSeek-R1 OpenAI+ECOD (16), OpenAI+LUNAR (6)
OpenAI-o1 OpenAI+DeepSVDD (11), OpenAI+iForest (7)
OpenAI-o3-mini BERT+DeepSVDD (10), OpenAI+ECOD (6)
@@ 5.3 Results, Insights, and Future Directions
The UMS scenario requires sophisticated reasoning. We select recent enhanced reasoning models, including OpenAI-o1-preview and OpenAI-o1 (OpenAI, 2024c), OpenAI-o3-mini (OpenAI, 2025), and DeepSeek-R1 (Guo et al., 2025).
LLM recommendations demonstrate strong potential. Figure 3 presents the model selection performance of four reasoning LLMs across five datasets, compared against two reference baselines: (i) the best result achieved by any baseline model, representing the performance upper bound; and (ii) the average performance of all baseline models, reflecting random model selection. In most cases, the AD performance of LLM-recommended models surpasses the average baseline and even approaches the best-performing model. These results highlight the strong potential of LLM-based reasoning to identify effective AD models using only public information, without reliance on historical performance or domain specialists.
LLMs exhibit inherent yet context-sensitive selection biases. The aggregated selection results in Table 3 highlight distinct model-selection preferences among the four LLMs. For example, OpenAI-o1-preview often recommends OpenAI+LUNAR, a consistently strong model in our benchmarks. In contrast, OpenAI-o3-mini prefers Bert+DeepSVDD, a generally weaker option. To investigate whether these biases arise from internal priors or context-specific information, we conducted a selection experiment without any dataset or model details (Table 4). Although each LLM still favors a distinct, limited set of models even without context, their specific preferences notably shift when introducing the context. This indicates that while LLMs possess intrinsic biases from their pretraining or tuning phases, their selections are also influenced by provided context information.
Context information improves selections, but justifications remain generic. When comparing model selection results with and without context (Table 3 - Table 4), we notice a clear shift in model recommendations that generally better align with benchmark results. Despite this improved selection accuracy, its explanations often remain generic and do not clearly link model selection to specific dataset characteristics. For example, in the “AG News” dataset, the OpenAI-o1-preview alternated between recommending “OpenAI + LUNAR” and “OpenAI + ECOD,” justifying choices with broad statements like “effective for high-dimensional data” or “parameter-free scalability.” Such non-specific rationales diminish interpretability and user trust, especially when understanding the rationale behind model choice is important.
Table 4: Selections by each LLM without any dataset or model context. Each LLM was queried five times.
LLMs Context-Free Picks (counts)
OpenAI-o1-preview OpenAI+LUNAR (3), OpenAI+VAE (2)
DeepSeek-R1 OpenAI+LUNAR (3), OpenAI+ECOD (2)
OpenAI-o1 OpenAI+VAE (4), OpenAI+LUNAR (1)
OpenAI-o3-mini OpenAI+VAE (4), OpenAI+DeepSVDD (1)
Future Direction 5: Refine Input Specificity and Alleviate Biases. Future work should explore how to provide more dataset-specific details and mitigate potential LLM biases. Ambiguous or incomplete input information may cause the LLM to favor well-known models or those frequently encountered during training. Ensuring detailed and balanced inputs, and exploring how inherent biases in LLMs affect recommendations, will be important steps to improve the fairness and reliability of LLM-based UMS Dai et al. (2024).
@@
6 Conclusion
In this work, we presented AD-LLM, the first comprehensive benchmark that integrates LLMs into three core aspects of anomaly detection in NLP: detection, data augmentation, and model selection.
Our results show that LLMs exhibit promising capabilities in zero-shot AD without task-specific training. LLM-generated synthetic data significantly boosted performance for models that learn flexible representations, while it may negatively impact models that rely on rigid geometric assumptions. Additionally, LLM-driven model selection frequently exceeded baseline performance, though explanations for these selections often lacked dataset-specific detail.
Future Directions
Future research should focus on improving contextual prompts to enhance zero-shot AD capabilities while considering the cost, developing methods to balance diversity and domain alignment in synthetic data generation, and increasing the specificity and interpretability of LLM-generated model selection justifications. Expanding the AD-LLM benchmark to include additional tasks and applications in different fields Huang et al. (2024b); Li et al. (2024b) also represents a valuable direction for broadening its impact.
@@ 6 Conclusion
AD-LLM explores the application of LLMs in enhancing AD through zero-shot detection, data augmentation, and model selection. These contributions have the potential to significantly improve real-world AD systems in critical areas such as healthcare, finance, and cybersecurity. By enabling robust, adaptable, and efficient solutions for AD tasks, this research empowers practitioners to deploy systems responsive to novel challenges while reducing reliance on labeled data and extensive domain expertise.
@@
This study adheres to ethical guidelines, emphasizing considerations around fairness, transparency, and privacy in developing and applying LLM-based AD systems. We emphasize the importance of evaluating and mitigating biases in LLM recommendations, ensuring that outputs are equitable and unbiased. Moreover, privacy is preserved by relying on public data and avoiding the collection of sensitive information. Also, note that we used ChatGPT exclusively to improve minor grammar in the final manuscript text.
@@
Despite promising results, several limitations remain. First, our evaluation is constrained to a narrow set of datasets with clear normal-anomaly distinctions, and our settings in AD and category descriptions in DA follow the structure of these datasets, limiting applicability to various domains with ambiguous anomaly definitions. Second, UMS depends on simplistic input data and matching mechanisms. Furthermore, biases in LLM recommendations, such as favoring well-documented or familiar models, need further investigation. Additionally, we do not explore few-shot learning or fine-tuning, which are widely adopted techniques for enhancing LLM performance and could offer valuable complementary insights for AD tasks.
@@ B.1 Prompt Details
  • Explanation and Implicit CoT. We require an explanation $r$ generated before the anomaly score $s$ for each inquiry as shown in Eq. (1). When generating the explanation, LLMs implicitly create the CoT in the background Liu et al. (2024d). This approach aligns with the auto-regressive nature of decoder-only LLMs, encouraging them to think carefully and logically before determining the anomaly score, thereby enhancing reliability.
In our experiments, we discovered that Llama 3.1 requires implicit CoT. Presenting the anomaly score $s$ before the explanation $r$ causes the Llama 3.1-based detector to crash and consistently outputs $s=0$. This issue does not impact GPT-4o and DeepSeek-V3. We attribute this to their significantly larger parameter count, which grants it a stronger resilience to prompt changes.
@@ B.3 Error Analysis
In our experiments, LLMs occasionally encountered errors. Llama 3.1 sometimes entered infinite loops, repeatedly generating the same sentence until reaching the token limit, as illustrated in the AG News example shown in Text Box B.3, where the repetitive nonsensical response is highlighted in italics. However, the underlying causes of such behavior remain underexplored. DeepSeek-V3 sometimes struggles to follow instructions, resulting in incorrect JSON formatting or failing to return any output. Additionally, it occasionally experiences internal server errors. For GPT-4o, outputs were sometimes restricted by OpenAI’s safety filters when addressing potentially harmful content, such as politically sensitive topics.
@@
Dataset “Normal Only” “Normal + Anomaly”
AG News 14 15
BBC News 3 1
IMDB Reviews 206 16
N24 News 138 252
SMS Spam 5 0
Table A5: Error count in DeepSeek-V3
We exclude these errors from our results but provide the error counts in both “Normal Only” and “Normal + Anomaly” settings for reference in Table A3 for Llama 3.1, Table A4 for GPT-4o, and Table A5 for DeepSeek-V3. Notably, the error counts vary between the two settings, suggesting that the triggers for errors, such as Llama’s infinite loop or GPT-4o’s safety filters, are sensitive to prompt variations. This occurs even though the prompts in both settings have similar semantic meanings.
B.4 Verbal Score
In the zero-shot AD task, we utilize LLM-generated verbal anomaly scores as a signal for detection. Verbalization methods are widely used because they offer an intuitive and straightforward estimation (Xia et al., 2025). However, LLMs can often be overconfident in their responses due to the influence of reinforcement learning from human feedback (RLHF) (Kadavath et al., 2022).
@@ C.1.1 Evaluation Protocol and Prompt Details
As discussed in §4.2, we set a scenario with limited training data $\mathcal{D}_{\text{small\_train}}=\left\{x_{1},x_{2},\dots,x_{m}\right\}$. Specifically, $\mathcal{D}_{\text{small\_train}}$ contains $v$ samples for each normal category $\mathcal{C}_{\text{normal}}^{j}\in\mathcal{C}_{\text{normal}}=\left\{\mathcal{C}_{\text{normal}}^{1},\dots,\mathcal{C}_{\text{normal}}^{k}\right\}$, where $k$ is the number of normal categories.
We employ a multi-step strategy with multiple rounds to mitigate repetitive outputs, token limit constraints, and difficulties in handling long contexts. The detailed pipeline is outlined below:
@@
    Keywords Generation. To ensure a consistent synthetic data distribution compared with the original training data, $t$ groups of keywords are generated for each normal category $\mathcal{C}_{\text{normal}}^{j}$. We construct the prompt $\mathcal{P}_{\text{keywords}}$ using a template $T_{\text{keywords}}(\cdot)$ as shown in Table  A11. This template utilizes {name} and {original_task} information from Table  A.1. The prompt $\mathcal{P}_{\text{keywords}}$ is processed by the LLM $f_{\text{LLM}}(\cdot)$ to produce $t\times k$ groups of keywords $\mathcal{K}=\left\{\mathcal{K}_{1},\mathcal{K}_{2},\dots,\mathcal{K}_{t\times k}\right\}$. Each keyword group $\mathcal{K}_{i}$ contains three keywords with increasing levels of granularity from coarse to fine.
@@
    Synthetic Sample Generation. We iterate the groups of keywords, constructing $\mathcal{P}_{\text{synth}}=\left\{\mathcal{P}_{\text{synth}}^{1},\mathcal{P}_{\text{synth}}^{2},\dots,\mathcal{P}_{\text{synth}}^{t\times k}\right\}$ using a template $T_{\text{synth}}(\cdot)$ as displayed in Table  A12. Each prompt $\mathcal{P}_{\text{synth}}^{j}$ is fed into the LLM $f_{\text{LLM}}(\cdot)$ to generate a corresponding synthetic sample $\hat{x}_{j}$. Finally, we obtain a synthetic dataset $\mathcal{D}_{\text{synth}}=\left\{\tilde{x}_{1},\tilde{x}_{2},\dots,\tilde{x}_{t\times k}\right\}$.
@@
$\mathcal{K}$ $=f_{\text{LLM}}\left(\mathcal{P}_{\text{keywords}}\right)=\left\{\mathcal{K}_{1},\dots,\mathcal{K}_{t\times k}\right\}$
$\mathcal{P}_{\text{synth}}$ =$=\left\{T_{\text{synth}}\left(\mathcal{K}_{1}\right),\dots,T_{\text{synth}}\left(\mathcal{K}_{t\times k}\right)\right\}$
$\mathcal{D}_{\text{synth}}$ =$=\left\{f_{\text{LLM}}\left(\mathcal{P}_{\text{synth}}^{1}\right),\dots,f_{\text{LLM}}\left(\mathcal{P}_{\text{synth}}^{t\times k}\right)\right\}$
@@ C.1.2 Experiments Details and Challenges
We set the number of samples from each normal category $\mathcal{C}_{\text{normal}}^{j}$ in the limited training set $\mathcal{D}_{\text{small\_train}}$ to $v=10$. Similarly, the number of synthetic samples generated for each normal category $\mathcal{C}_{\text{normal}}^{j}$ in the synthetic set $\mathcal{D}_{\text{synth}}$ is set to $t=50$ for the “AG New”, “BBC News”, “IMDB Reviews”, and “SMS Spam” datasets. For the “N24 News” dataset, we set $v=3$ and $t=30$ to account for its numerous normal categories.
@@
We use GPT-4o for synthetic data generation. We observed that increasing $t$ occasionally causes them to terminate the keyword generation process before reaching the token limit. A similar issue occurs with Llama 3.1, even for smaller values of $t$. As a result, Llama 3.1 is excluded from this task. We presume these issues stem from the inherent challenges LLMs face in processing long contexts. We also exclude DeepSeek-V3 due to its unsatisfactory results.
Table A6: Performance comparison of AD baselines with and without LLM-generated synthetic data across five datasets. We also show the average performance $\pm$ its standard deviation over five datasets. The better results for each detector are highlighted in bold. The performance may vary due to the embedding changes.
Training Set AG News BBC News IMDB Reviews N24 News SMS Spam Average Performance
AUROC $\uparrow$ AUPRC $\uparrow$ AUROC $\uparrow$ AUPRC $\uparrow$ AUROC $\uparrow$ AUPRC $\uparrow$ AUROC $\uparrow$ AUPRC $\uparrow$ AUROC $\uparrow$ AUPRC $\uparrow$ AUROC $\uparrow$ AUPRC $\uparrow$
OpenAI + AE
without $\mathcal{D}_{\text{synth}}$ 0.5054 0.1189 0.6016 0.1309 0.5014 0.1665 0.7119 0.1681 0.5000 0.1020 0.5641$\pm$0.0834 0.1373$\pm$0.0262
with $\mathcal{D}_{\text{synth}}$ 0.8097 0.3290 0.8434 0.3936 0.8097 0.3290 0.8097 0.3290 0.4341 0.0852 0.7413$\pm$0.1542 0.2932$\pm$0.1069
OpenAI + DeepSVDD
without $\mathcal{D}_{\text{synth}}$ 0.5171 0.1237 0.6127 0.1415 0.5667 0.1969 0.6278 0.1511 0.6398 0.1479 0.5928$\pm$0.0453 0.1522$\pm$0.0243
with $\mathcal{D}_{\text{synth}}$ 0.5554 0.1365 0.5867 0.1267 0.5554 0.1365 0.5554 0.1365 0.3086 0.0681 0.5123$\pm$0.1026 0.1209$\pm$0.0267
OpenAI + ECOD
without $\mathcal{D}_{\text{synth}}$ 0.5014 0.1180 0.5623 0.1208 0.5000 0.1661 0.6202 0.1311 0.4078 0.0789 0.5183$\pm$0.0709 0.1230$\pm$0.0279
with $\mathcal{D}_{\text{synth}}$ 0.6709 0.1954 0.7660 0.3210 0.6709 0.1954 0.6709 0.1954 0.3351 0.0708 0.6228$\pm$0.1485 0.1956$\pm$0.0791
OpenAI + IForest
without $\mathcal{D}_{\text{synth}}$ 0.6120 0.1620 0.7102 0.1903 0.5788 0.1947 0.5331 0.1010 0.6386 0.1467 0.6145$\pm$0.0594 0.1589$\pm$0.0340
with $\mathcal{D}_{\text{synth}}$ 0.6759 0.2159 0.6655 0.2107 0.6759 0.2159 0.6759 0.2159 0.2700 0.0649 0.5926$\pm$0.1614 0.1847$\pm$0.0599
OpenAI + LOF
without $\mathcal{D}_{\text{synth}}$ 0.6404 0.1661 0.7128 0.2565 0.6759 0.2485 0.7179 0.2061 0.7582 0.2445 0.7010$\pm$0.0400 0.2243$\pm$0.0339
with $\mathcal{D}_{\text{synth}}$ 0.5469 0.1411 0.6513 0.2075 0.5469 0.1411 0.5469 0.1411 0.8150 0.2602 0.6214$\pm$0.1049 0.1782$\pm$0.0484
OpenAI + SO_GAAL
without $\mathcal{D}_{\text{synth}}$ 0.5657 0.1324 0.3240 0.0770 0.5388 0.1659 0.3351 0.0654 0.3953 0.0823 0.4318$\pm$0.1017 0.1046$\pm$0.0383
with $\mathcal{D}_{\text{synth}}$ 0.4461 0.0976 0.2787 0.0703 0.4461 0.0976 0.4461 0.0976 0.0698 0.0637 0.3374$\pm$0.1487 0.0854$\pm$0.0151
OpenAI + LUNAR
without $\mathcal{D}_{\text{synth}}$ 0.6527 0.2035 0.8554 0.4670 0.6546 0.2315 0.7879 0.2473 0.1506 0.0573 0.6202$\pm$0.2475 0.2413$\pm$0.1314
with $\mathcal{D}_{\text{synth}}$ 0.8651 0.4228 0.9330 0.7332 0.8651 0.4228 0.8651 0.4228 0.1375 0.0568 0.7332$\pm$0.2990 0.4117$\pm$0.2143
OpenAI + VAE
without $\mathcal{D}_{\text{synth}}$ 0.6857 0.1842 0.7143 0.1816 0.5031 0.1670 0.6932 0.1698 0.5000 0.1020 0.6193$\pm$0.0966 0.1609$\pm$0.0302
with $\mathcal{D}_{\text{synth}}$ 0.7905 0.3654 0.7674 0.2654 0.7905 0.3654 0.7905 0.3654 0.0696 0.0545 0.6417$\pm$0.2862 0.2832$\pm$0.1207
We repeat the generation of keywords four times, with different temperatures $[1.0,0.9,0.8,0.7]$ and different seeds $[42,43,44,45]$. To further avoid repetition, we add additional sentences to the end of the prompts, including:
  • “This is the first time you do this task, good luck!"
  • “You’ve completed this task before, and you’re improving at it."
  • “After doing this task twice, you have a better understanding of it."
  • “You have done this task three times, you are now an expert at it."
Table A7: Complete model selection results across five datasets. We display the average AUROC and AUPRC of models recommended by querying each reasoning LLM five times (duplicates allowed). "Best Performance" marks the highest performance achieved by any baseline model for each dataset, while "Average Performance" denotes the mean performance across all baseline models.
Settings AG News BBC News IMDB Reviews N24 News SMS Spam
OpenAI-o1 0.7132 0.3199 0.6798 0.2831 0.6563 0.3278 0.7091 0.2419 0.3647 0.0752
OpenAI-o1-preview 0.8908 0.6193 0.6992 0.2214 0.6652 0.2787 0.7706 0.3422 0.5774 0.1220
OpenAI-o3-mini 0.6455 0.2401 0.7329 0.3132 0.5358 0.2521 0.4870 0.0952 0.5758 0.1169
DeepSeek-R1 0.8273 0.4744 0.7224 0.2424 0.7009 0.3976 0.7733 0.3113 0.5090 0.1022
We carefully examine and remove duplicate keyword groups. Out of 200 generations, there are typically fewer than 5 repeated groups, with a maximum of 15. It shows that our method is effective.